OpenStudy (anonymous):

FAN AND MEDAL ALGEBRA 1 QUESTIONS

OpenStudy (anonymous):

What are they?

OpenStudy (anonymous):

Solve for x: −4|x + 5| = −16 x = 21/4 , x = −11/4 x = −1, x = 9 x = −1, x = −9 No solution

OpenStudy (anonymous):

Are those parentheses?

OpenStudy (saifoo.khan):